# A Weighted Loss Approach to Robust Federated Learning under Data Heterogeneity

### v2

@@ 8.6 Impact of Batch Size & Momentum on WoLA (RQ6)
9 Conclusion and Future Work
We show that label skew heterogeneity can be partly addressed with WoLA, a weighted loss where weights are set according to each worker’s local label distribution and a shared training objective. WoLA is a simple, model-agnostic method that requires no hyperparameter tuning, no server-side dataset, no knowledge of the number of Byzantine workers, and adds no computational overhead. It is easy to implement and can be combined with other robustness strategies, such as pre-aggregators and robust aggregators.
We provide theoretical insights into WoLA’s behavior, supported by empirical evaluation. Extensive experiments demonstrate WoLA’s efficiency and its superiority over state-of-the-art strategies—both on average and in worst-case scenarios (i.e., under the strongest attack for each method)—especially in adversarial settings involving both Byzantine behavior and heterogeneity. Even better results are achieved when WoLA is combined with NNM. We also show that WoLA significantly reduces gradient dissimilarity among honest workers, which is key to its robustness.
For future work, WoLA’s robustness can be improved and its scope extended—for instance, by refining the weighting scheme, such as introducing thresholds or adapting it dynamically at the mini-batch scale. The impact of training objective attacks can also be explored more thoroughly, along with the design of stronger defenses against them. Finally, developing a decentralized version of WoLA also seems to be a relevant direction.
Appendix A Theoretical Results
In this section, we define the random counterparts of the local updates and provide all proofs of the theoretical results given in Section 6.2.
A.1 Gradients Modeling
In this subsection, we define the random counterparts of $\nabla\mathcal{L}_{i}$ and $\nabla\mathcal{WL}_{i}$. Let the global dataset be $\mathcal{D}=\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{N},y_{N})\}$ with $(x_{i},y_{i})\in\mathcal{X}\times[C]$, and let $z_{1},z_{2},\ldots,z_{N}$ be integers in $\mathcal{H}$ such that $z_{k}=i$ if $(x_{k},y_{k})\in\mathcal{D}_{i}$.
Since $N_{i}=\sum_{l=1}^{N}\mathbbm{1}(z_{l}=i)$, the following holds for $\nabla\mathcal{L}_{i}$:
$=$ $\frac{N}{N_{i}}\frac{\sum_{k=1}^{N}\mathbbm{1}(z_{k}=i)\ell\left(y_{k},\Phi\left(x_{k}\right)\right)}{N}$
$=$ $\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(z_{l}=i)}{N}\Bigg{)}^{-1}\frac{\sum_{k=1}^{N}\mathbbm{1}(z_{k}=i)\nabla\ell(y_{k},\Phi(x_{k}))}{N}$
Let $L_{i}$ be the random counterpart of $\nabla\mathcal{L}_{i}$. Replacing realizations with random variables (as defined in Subsection 6.1), $L_{i}$ is given by:
$L_{i}=$ $\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$ (5)
$\quad\quad\frac{\sum_{k=1}^{N}\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}$
We now turn to $\nabla\mathcal{WL}_{i}$:
$\nabla\mathcal{WL}_{i}=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\frac{q^{y}}{p^{y}_{i}}\nabla\ell\Big{(}y,\Phi(x)\Big{)}$
$=$ $\frac{1}{N_{i}}\sum_{k=1}^{N}\mathbbm{1}(z_{k}=i)\frac{q^{y_{k}}}{p^{y_{k}}_{i}}\nabla\ell\Big{(}y_{k},\Phi(x_{k})\Big{)}$
$=$ $\frac{1}{N_{i}}\sum_{k=1}^{N}\mathbbm{1}(z_{k}=i)\sum_{c=1}^{C}\mathbbm{1}(y_{k}=c)\frac{q^{c}}{N_{i}^{c}/N_{i}}\nabla\ell\Big{(}y_{k},\Phi(x_{k})\Big{)}$
$=$ $\sum_{c=1}^{C}q^{c}\frac{N}{N_{i}^{c}}\frac{\sum_{k=1}^{N}\mathbbm{1}(y_{k}=c)\mathbbm{1}(z_{k}=i)\nabla\ell\Big{(}y_{k},\Phi(x_{k})\Big{)}}{N}$
Since $N^{c}_{i}=\sum_{l=1}^{N}\mathbbm{1}(y_{l}=c)\mathbbm{1}(z_{l}=i)$, we have:
$\nabla\mathcal{WL}_{i}=$ $\sum_{c=1}^{C}q^{c}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(y_{l}=c)\mathbbm{1}(z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbbm{1}(y_{k}=c)\mathbbm{1}(z_{k}=i)\nabla\ell(y_{k},\Phi(x_{k}))}{N}\Bigg{)}$
Let $W_{i}$ be the random counterpart of $\nabla\mathcal{WL}_{i}$. Replacing realizations with random variables (as defined in Subsection 6.1), $W_{i}$ is given by:
$W_{i}=$ $\sum_{c=1}^{C}q^{c}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Y_{l}=c)\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$ (6)
$\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbbm{1}(Y_{k}=c)\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
A.2 Lemmas
Lemma 2
Proof.
$\mathbb{E}\Big{[}f(X)\Big{|}Y=y,Z=z\Big{]}=\frac{\mathbb{E}\Big{[}\mathbbm{1}(Y=y)\mathbbm{1}(Z=z)f(X)\Big{]}}{\mathbb{P}(Y=y,Z=z)}$
$=$ $\mathbb{E}\Big{[}\mathbbm{1}(Z=z)f(X)\Big{|}Y=y\Big{]}\frac{\mathbb{P}(Y=y)}{\mathbb{P}(Y=y,Z=z)}$
$=$ $\mathbb{E}\Big{[}\mathbbm{1}(Z=z)\Big{|}Y=y\Big{]}\mathbb{E}\Big{[}f(X)\Big{|}Y=y\Big{]}\frac{\mathbb{P}(Y=y)}{\mathbb{P}(Y=y,Z=z)}$
$=$ $\frac{\mathbb{E}\Big{[}\mathbbm{1}(Z=z)\mathbbm{1}(Y=y)\Big{]}}{\mathbb{P}(Y=y)}\mathbb{E}\Big{[}f(X)\Big{|}Y=y\Big{]}\frac{\mathbb{P}(Y=y)}{\mathbb{P}(Y=y,Z=z)}$
$=$ $\mathbb{E}\Big{[}f(X)\Big{|}Y=y\Big{]}$
Lemma 3
$\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Y_{l}=c)\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbbm{1}(Y_{k}=c)\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
[…]
Proof.
$\lim_{N\to\infty}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Y_{l}=c)\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\quad\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbbm{1}(Y_{k}=c)\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
$=\Bigg{(}\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbbm{1}(Y_{l}=c)\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\quad\quad\Bigg{(}\lim_{N\to\infty}\frac{\sum_{k=1}^{N}\mathbbm{1}(Y_{k}=c)\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
$\mathbb{E}\Big{[}\mathbbm{1}(Y=c)\mathbbm{1}(Z=i)\Big{]}^{-1}$
$\quad\quad\quad\quad\mathbb{E}\Big{[}\mathbbm{1}(Y=c)\mathbbm{1}(Z=i)\nabla\ell(Y,\Phi(X))\Big{]}$
$=$ $\frac{\mathbb{E}\Big{[}\mathbbm{1}(Y=c)\mathbbm{1}(Z=i)\nabla\ell(Y,\Phi(X))\Big{]}}{\mathbb{P}(Y=c,Z=i)}$
$=$ $\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c,Z=i\Big{]}$
According to Lemma 2:
A.3 Proof of Proposition 1
We begin by calculating the limit of $L_{i}$. By definition, we have:
$L_{i}=\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}\frac{\sum_{k=1}^{N}\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}$
By the Law of Large Numbers, it holds:
[…]
$\frac{\sum_{k=1}^{N}\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}$
[…]
$\lim_{N\to\infty}L_{i}=\lim_{N\to\infty}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\quad\quad\frac{\sum_{k=1}^{N}\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}$
$=$ $\Bigg{(}\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\quad\quad\Bigg{(}\lim_{N\to\infty}\frac{\sum_{k=1}^{N}\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
[…]
Similarly, for $L$:
$\lim_{N\to\infty}L=$ $\lim_{N\to\infty}\sum_{i\in\mathcal{H}}L_{i}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}$
$=$ $\sum_{i\in\mathcal{H}}\lim_{N\to\infty}L_{i}\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}$
[…]
By the Law of total expectation, it holds:
[…]
A.4 Proof of Proposition 2
$\lim_{N\to\infty}W_{i}$
$=$ $\lim_{N\to\infty}\sum_{c=1}^{C}q^{c}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Y_{l}=c)\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\quad\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbbm{1}(Y_{k}=c)\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
$=$ $\sum_{c=1}^{C}q^{c}\lim_{N\to\infty}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbbm{1}(Y_{l}=c)\mathbbm{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\quad\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbbm{1}(Y_{k}=c)\mathbbm{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
[…]
Similarly, for $W$:
$\lim_{N\to\infty}W=$ $\lim_{N\to\infty}\sum_{i\in\mathcal{H}}W_{i}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}$
$=$ $\sum_{i\in\mathcal{H}}\lim_{N\to\infty}W_{i}\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbbm{1}(Z_{l}=i)}{N}$
By the Law of Large Numbers, it holds:
[…]
Using above results, it concludes the lemma:
[…]
A.5 Proof of Gradient Dissimilarity Vanishing
[…]
A.6 Proof of Proposition 3
According to Proposition 2, it holds:
[…]
$=\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{]}$
A.7 Worst Training Objective Attack
@@
Appendix B Experiment
B.1 Robust Distributed Stochastic Heavy Ball Algorithm

### v3

@@ Appendix D Additional Research Questions
Appendix E Training Objective Attack
Our experiments include scenarios where workers share local label distributions with the server, while Byzantine machines manipulate the training objective by sending malicious distributions. We describe this attack setting below.
WoLA Scenario
Similar to how Byzantine workers can send arbitrary gradient updates with full knowledge of honest gradients, we assume they can also send arbitrary label distributions with full knowledge of the honest workers’ local label distributions. Rather than enumerating all possible attacks, we focus on a worst-case attack: Byzantine machines submit a zero vector except for a single entry set to $1$, corresponding to the minority class in the global dataset (proofs are provided in the next subsection).
To defend against such attacks, we adopt the geometric median as the aggregation method. This choice offers two main advantages: (i) Unlike methods such as Trimmed Mean or Krum, which exclude outliers, the geometric median treats all updates equally. This is crucial since outlier distributions are not necessarily malicious; excluding them risks discarding honest data and harming performance. (ii) When aggregating probability distributions, the geometric median produces a valid probability distribution—unlike the coordinate-wise median—thus preserving the training objective’s expected form.
Our aim is to show that even a worst-case attack has limited impact on model performance when countered by a simple defense. We denote our method under this attack scenario as WoLA.
Proof that WoLA is a Worst-Case Attack
@@ Appendix E Training Objective Attack
Appendix F Theoretical Proofs
In this section, we define the random counterparts of the local updates and provide all proofs of the theoretical results given in Section WoLA Properties.
Gradients Modeling
In this subsection, we define the random counterparts of $\nabla\mathcal{L}_{i}$ and $\nabla\mathcal{W}\!\mathcal{L}_{i}$. Let the global dataset be $\mathcal{D}=\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{N},y_{N})\}$ with $(x_{i},y_{i})\in\mathcal{X}\times[C]$, and let $z_{1},z_{2},\ldots,z_{N}$ be integers in $\mathcal{H}$ such that $z_{k}=i$ if $(x_{k},y_{k})\in\mathcal{D}_{i}$.
Since $N_{i}=\sum_{l=1}^{N}\mathbf{1}(z_{l}=i)$, the following holds for $\nabla\mathcal{L}_{i}$:
$=$ $\frac{N}{N_{i}}\frac{\sum_{k=1}^{N}\mathbf{1}(z_{k}=i)\ell\left(y_{k},\Phi\left(x_{k}\right)\right)}{N}$
$=$ $\Bigg{(}\frac{\sum_{l=1}^{N}\mathbf{1}(z_{l}=i)}{N}\Bigg{)}^{-1}\frac{\sum_{k=1}^{N}\mathbf{1}(z_{k}=i)\nabla\ell(y_{k},\Phi(x_{k}))}{N}$
Let $L_{i}$ be the random counterpart of $\nabla\mathcal{L}_{i}$. Replacing realizations with random variables (as defined in Subsection Theoretical Framework), $L_{i}$ is given by:
$L_{i}=$ $\Bigg{(}\frac{\sum_{l=1}^{N}\mathbf{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$ (2)
$\quad\quad\frac{\sum_{k=1}^{N}\mathbf{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}$
We now turn to $\nabla\mathcal{W}\!\mathcal{L}_{i}$:
$\nabla\mathcal{W}\!\mathcal{L}_{i}=\frac{1}{N_{i}}\sum_{(x,y)\in\mathcal{D}_{i}}\frac{q^{y}}{p^{y}_{i}}\nabla\ell\Big{(}y,\Phi(x)\Big{)}$
$=$ $\frac{1}{N_{i}}\sum_{k=1}^{N}\mathbf{1}(z_{k}=i)\frac{q^{y_{k}}}{p^{y_{k}}_{i}}\nabla\ell\Big{(}y_{k},\Phi(x_{k})\Big{)}$
$=$ $\frac{1}{N_{i}}\sum_{k=1}^{N}\mathbf{1}(z_{k}=i)\sum_{c=1}^{C}\mathbf{1}(y_{k}=c)\frac{q^{c}}{N_{i}^{c}/N_{i}}\nabla\ell\Big{(}y_{k},\Phi(x_{k})\Big{)}$
$=$ $\sum_{c=1}^{C}q^{c}\frac{N}{N_{i}^{c}}\frac{\sum_{k=1}^{N}\mathbf{1}(y_{k}=c)\mathbf{1}(z_{k}=i)\nabla\ell\Big{(}y_{k},\Phi(x_{k})\Big{)}}{N}$
Since $N^{c}_{i}=\sum_{l=1}^{N}\mathbf{1}(y_{l}=c)\mathbf{1}(z_{l}=i)$, we have:
$\nabla\mathcal{W}\!\mathcal{L}_{i}=$ $\sum_{c=1}^{C}q^{c}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbf{1}(y_{l}=c)\mathbf{1}(z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbf{1}(y_{k}=c)\mathbf{1}(z_{k}=i)\nabla\ell(y_{k},\Phi(x_{k}))}{N}\Bigg{)}$
Let $W_{i}$ be the random counterpart of $\nabla\mathcal{W}\!\mathcal{L}_{i}$. Replacing realizations with random variables (as defined in Subsection Theoretical Framework), $W_{i}$ is given by:
$W_{i}=$ $\sum_{c=1}^{C}q^{c}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbf{1}(Y_{l}=c)\mathbf{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$ (3)
$\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbf{1}(Y_{k}=c)\mathbf{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
Lemmas
Lemma 1
Proof.
$\mathbb{E}\Big{[}f(X)\Big{|}Y=y,Z=z\Big{]}=\frac{\mathbb{E}\Big{[}\mathbf{1}(Y=y)\mathbf{1}(Z=z)f(X)\Big{]}}{\mathbb{P}(Y=y,Z=z)}$
$=$ $\mathbb{E}\Big{[}\mathbf{1}(Z=z)f(X)\Big{|}Y=y\Big{]}\frac{\mathbb{P}(Y=y)}{\mathbb{P}(Y=y,Z=z)}$
$=$ $\mathbb{E}\Big{[}\mathbf{1}(Z=z)\Big{|}Y=y\Big{]}\mathbb{E}\Big{[}f(X)\Big{|}Y=y\Big{]}\frac{\mathbb{P}(Y=y)}{\mathbb{P}(Y=y,Z=z)}$
$=$ $\frac{\mathbb{E}\Big{[}\mathbf{1}(Z=z)\mathbf{1}(Y=y)\Big{]}}{\mathbb{P}(Y=y)}\mathbb{E}\Big{[}f(X)\Big{|}Y=y\Big{]}\frac{\mathbb{P}(Y=y)}{\mathbb{P}(Y=y,Z=z)}$
$=$ $\mathbb{E}\Big{[}f(X)\Big{|}Y=y\Big{]}$
Lemma 2
$\Bigg{(}\frac{\sum_{l=1}^{N}\mathbf{1}(Y_{l}=c)\mathbf{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbf{1}(Y_{k}=c)\mathbf{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
$\quad\quad\quad\quad\underset{N\rightarrow\infty}{\overset{\text{a.s.}}{\longrightarrow}}\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c\Big{]}$
Proof.
$\lim_{N\to\infty}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbf{1}(Y_{l}=c)\mathbf{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\quad\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbf{1}(Y_{k}=c)\mathbf{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
$=\Bigg{(}\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbf{1}(Y_{l}=c)\mathbf{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\quad\quad\Bigg{(}\lim_{N\to\infty}\frac{\sum_{k=1}^{N}\mathbf{1}(Y_{k}=c)\mathbf{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
$\mathbb{E}\Big{[}\mathbf{1}(Y=c)\mathbf{1}(Z=i)\Big{]}^{-1}$
$\quad\quad\quad\quad\mathbb{E}\Big{[}\mathbf{1}(Y=c)\mathbf{1}(Z=i)\nabla\ell(Y,\Phi(X))\Big{]}$
$=$ $\frac{\mathbb{E}\Big{[}\mathbf{1}(Y=c)\mathbf{1}(Z=i)\nabla\ell(Y,\Phi(X))\Big{]}}{\mathbb{P}(Y=c,Z=i)}$
$=$ $\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c,Z=i\Big{]}$
According to Lemma 1:
Proof of Proposition 1
We begin by calculating the limit of $L_{i}$. By definition, we have:
$L_{i}=\Bigg{(}\frac{\sum_{l=1}^{N}\mathbf{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}\frac{\sum_{k=1}^{N}\mathbf{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}$
By the Law of Large Numbers, it holds:
$\frac{\sum_{l=1}^{N}\mathbf{1}(Z_{l}=i)}{N}\underset{N\rightarrow\infty}{\overset{\text{a.s.}}{\longrightarrow}}\mathbb{E}\Big{[}\mathbf{1}(Z=i)\Big{]}=\mathbb{P}(Z=i),$
$\frac{\sum_{k=1}^{N}\mathbf{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}$
$\quad\quad\quad\quad\underset{N\rightarrow\infty}{\overset{\text{a.s.}}{\longrightarrow}}\mathbb{E}[\mathbf{1}(Z=i)\nabla\ell(Y,\Phi(X))].$
$\lim_{N\to\infty}L_{i}=\lim_{N\to\infty}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbf{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\quad\quad\frac{\sum_{k=1}^{N}\mathbf{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}$
$=$ $\Bigg{(}\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbf{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\quad\quad\Bigg{(}\lim_{N\to\infty}\frac{\sum_{k=1}^{N}\mathbf{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
$\overset{\text{a.s.}}{=}$ $\frac{\mathbb{E}[\mathbf{1}(Z=i)\nabla\ell(Y,\Phi(X))]}{\mathbb{P}(Z=i)}=\mathbb{E}[\nabla\ell(Y,\Phi(X))|Z=i]$
Similarly, for $L$:
$\lim_{N\to\infty}L=$ $\lim_{N\to\infty}\sum_{i\in\mathcal{H}}L_{i}\frac{\sum_{l=1}^{N}\mathbf{1}(Z_{l}=i)}{N}$
$=$ $\sum_{i\in\mathcal{H}}\lim_{N\to\infty}L_{i}\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbf{1}(Z_{l}=i)}{N}$
$\overset{\text{a.s.}}{=}$ $\sum_{i\in\mathcal{H}}\mathbb{E}[\nabla\ell(Y,\Phi(X))|Z=i]\mathbb{P}(Z=i)$
By the Law of total expectation, it holds:
$\lim_{N\to\infty}L\overset{\text{a.s.}}{=}\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{]}$
Proof of Proposition 2
$\lim_{N\to\infty}W_{i}$
$=$ $\lim_{N\to\infty}\sum_{c=1}^{C}q^{c}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbf{1}(Y_{l}=c)\mathbf{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\quad\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbf{1}(Y_{k}=c)\mathbf{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
$=$ $\sum_{c=1}^{C}q^{c}\lim_{N\to\infty}\Bigg{(}\frac{\sum_{l=1}^{N}\mathbf{1}(Y_{l}=c)\mathbf{1}(Z_{l}=i)}{N}\Bigg{)}^{-1}$
$\quad\quad\quad\quad\Bigg{(}\frac{\sum_{k=1}^{N}\mathbf{1}(Y_{k}=c)\mathbf{1}(Z_{k}=i)\nabla\ell(Y_{k},\Phi(X_{k}))}{N}\Bigg{)}$
$\lim_{N\to\infty}W_{i}\overset{\text{a.s.}}{=}\sum_{c=1}^{C}q^{c}\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c\Big{]}$
Similarly, for $W$:
$\lim_{N\to\infty}W=$ $\lim_{N\to\infty}\sum_{i\in\mathcal{H}}W_{i}\frac{\sum_{l=1}^{N}\mathbf{1}(Z_{l}=i)}{N}$
$=$ $\sum_{i\in\mathcal{H}}\lim_{N\to\infty}W_{i}\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbf{1}(Z_{l}=i)}{N}$
By the Law of Large Numbers, it holds:
$\lim_{N\to\infty}\frac{\sum_{l=1}^{N}\mathbf{1}(Z_{l}=i)}{N}\overset{\text{a.s.}}{=}\mathbb{E}\Big{[}\mathbf{1}(Z=i)\Big{]}=\mathbb{P}(Z=i),$
Using above results, it concludes the lemma:
$\lim_{N\to\infty}W\overset{\text{a.s.}}{=}\sum_{i\in\mathcal{H}}\sum_{c=1}^{C}q^{c}\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c\Big{]}\mathbb{P}(Z=i)$
Proof of Gradient Dissimilarity Vanishing
$=$ $\frac{1}{\#\mathcal{H}}\sum_{i\in\mathcal{H}}\left\|\lim_{N\to\infty}W_{i}-\lim_{N\to\infty}W\right\|^{2}\overset{\text{a.s.}}{=}0$
Proof of Proposition 3
According to Proposition 2, it holds:
$W\underset{N\rightarrow\infty}{\overset{\text{a.s.}}{\longrightarrow}}\sum_{c=1}^{C}q^{c}\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c\Big{]}$
$W\underset{N\rightarrow\infty}{\overset{\text{a.s.}}{\longrightarrow}}$ $\sum_{c=1}^{C}\mathbb{P}(Y=c)\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{|}Y=c\Big{]}$
$=\mathbb{E}\Big{[}\nabla\ell(Y,\Phi(X))\Big{]}$
Appendix G Experimental Details
Robust Distributed Stochastic Heavy Ball Algorithm